OpenStudy (anonymous):

please help me solve this equation using trigonometric integrals. ʃcsc^6xdx

OpenStudy (wasiqss):

w8

OpenStudy (anonymous):

This may be of assistance -> http://en.wikipedia.org/wiki/Integration_by_reduction_formulae

OpenStudy (unklerhaukus):

\[\int sin^{-6}(x)dx\]

OpenStudy (wasiqss):

\[\int\limits_{?}^{?}i-\cos^3(2x)\]

OpenStudy (wasiqss):

sorry not i , its 1

OpenStudy (wasiqss):

its easier now

OpenStudy (wasiqss):

integral of 1 is x, so now we only have to integrate cos^3(2x)

OpenStudy (wasiqss):

now we can apply reduction formula

OpenStudy (wasiqss):

answer is \[x+\sin2xcos^2(2x)+4\sin2x\]

OpenStudy (anonymous):

how did ʃcsc^6xdx become ʃcos^3(2x)?

OpenStudy (wasiqss):

it became 1+cos^3(2x)

OpenStudy (wasiqss):

by applying identity i did a minor mistake it shud b 1/2+1/2cos^3(2x)

OpenStudy (wasiqss):

by applying identity cos2x=1-2sin^2x

OpenStudy (wasiqss):

then i applied reduction formula to cos^3(2x)

OpenStudy (anonymous):

thank you :)

OpenStudy (wasiqss):

u can check the reduction formula fsm jus posted above

OpenStudy (anonymous):

okay C: